OpenStudy (anonymous):

I need help with proving the identity csc(x) - cos(x) cot(x) = sin(x)? Will give medal & fan. All I need for this is the reasoning. I have the proof step-by-step, but I need the reasons for these steps. I'll be posting the steps in the comments.

OpenStudy (anonymous):

Step 1 csc(x) - cos(x) cot(x) = (1/sin(x)) - cos(x)(cos(x)/sin(x))

OpenStudy (hba):

Cosec(x)=1/Sinx and Cotx=Cosx/Sinx

OpenStudy (anonymous):

Step 2 csc(x) - cos(x) cot(x) = (1/sin(x)) - (cos^2 (x) / sin(x))

OpenStudy (hba):

cosx*cosx=cos^2(x)

OpenStudy (anonymous):

Step 3 csc(x) - cos(x) cot(x) = (1 - cos^2 (x)) / sin(x)

OpenStudy (hba):

Lcm on both side.